OpenStudy (anonymous):

\[^{4}\sqrt{5x-9}=5\]

OpenStudy (amistre64):

5x-9=25.25 5x = 25.25 + 9 x = 5.25 + 9 x = 125+9 = 134

OpenStudy (amistre64):

ack..forgot somethin

OpenStudy (amistre64):

x = 125 + 9/5 lol

OpenStudy (amistre64):

125 + 1' 4/5 126' 4/5

OpenStudy (anonymous):

the = sign is supposed to be < or = too 5

OpenStudy (anonymous):

how are you getting 25.25??

OpenStudy (anonymous):

im sorry

OpenStudy (yuki):

amistre... lol

OpenStudy (amistre64):

25.25 is lazy muliplication lol

OpenStudy (yuki):

littleice, this will be the last problem that I can help since I need to go soon ok?

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

5x-9 >= 5.5.5.5 like that?

OpenStudy (yuki):

\[^4\sqrt{5x-9} = 5\] the thing is that the left side is not a sqrt, its a 4th rt so in order to get rid of that you need to take both sides to the "4th" power

OpenStudy (yuki):

so 5x-9 = 625 5x = 634 thus x=634/5 okay?

OpenStudy (amistre64):

5x -9 <= 5.5.5.5 5x <= 5.5.5.5 +9 x = <5.5.5.+ 9/5 x<= 125 + 1'4/5 x <= 126' 4/5 :)

OpenStudy (amistre64):

x is between 9/5 and that upper limit

OpenStudy (anonymous):

... amistre, x is an exact value.

OpenStudy (amistre64):

...i undertstand your confusion; but ice modified it to include the <=

OpenStudy (anonymous):

oh he did? XD

OpenStudy (anonymous):

answer is 1.8 is < or = x and is < than or = to 5

OpenStudy (anonymous):

sry XD